### v1

# Estimating treatment effects with competing intercurrent events in randomized controlled trials

Abstract
The analysis of randomized controlled trials is often complicated by intercurrent events—events that occur after treatment initiation and may impact outcome assessment. These events may lead to patients discontinuing their assigned treatment or dropping out of the trial entirely. In an analysis of data from two recent immunology trials, we categorize intercurrent events into two broad types: those unrelated to treatment (e.g., withdrawal from the study due to external factors like pandemics or relocation) and those related to treatment (e.g., adverse events or lack of efficacy). We adopt distinct strategies to handle each type, aiming to target a clinically more relevant estimand. For treatment-related intercurrent events, they often meaningfully describe the patient’s outcome, we employ a composite variable strategy, where we attribute an outcome value that reflects the lack of treatment success. For treatment-unrelated intercurrent events, we adopt a hypothetical strategy that assumes these event times are conditionally independent of the outcome, given treatment and covariates, and envisions a scenario in which the intercurrent events do not occur. We establish the nonparametric identification and semiparametric estimation theory for the causal estimand and introduce doubly robust estimators. We illustrate our methods through the re-analysis of two randomized trials on baricitinib for Systemic Lupus Erythematosus. We classify intercurrent events, apply four estimators, and compare our approach with common ad-hoc methods, highlighting the robustness and practical implications of our framework.
Keywords: Causal inference; Censoring; Estimand; Post-treatment variable; Potential outcome
1 Intercurrent events in randomized controlled trials
Randomized controlled trials (RCTs) are considered the gold standard for evaluating treatment efficacy. However, after treatment initiation, various events—referred to as intercurrent events (ICEs)—can arise, impacting the interpretation or availability of the outcome of interest, and posing significant challenges to the analysis of RCTs. These ICEs can include treatment discontinuation due to adverse events, lack of efficacy, or study withdrawal due to patient relocation. Addressing these ICEs carefully is crucial to ensure the validity and reliability of the causal conclusions drawn from RCTs.
Given the complexities introduced by ICEs, clear guidance on how to handle them is crucial for drawing valid causal conclusions. The International Conference on Harmonization (ICH) has established global standards for drug development to align practices among pharmaceutical companies and regulatory agencies. Recognizing the importance of well-defined estimands in the presence of ICEs, the ICH introduced the E9 (R1) Addendum (ICH E9 , 2019, R1), a key resource that provides updated statistical principles to support clear and consistent approaches to analyzing and interpreting RCT results. The guidance has become essential for companies in drug research and development, and various strategies have been proposed and utilized to address ICEs following the guidance (Qu et al.,, 2021; Han and Zhou,, 2023).
One straightforward approach to address the complications from the ICEs is the intention-to-treat (ITT) analysis. The ITT strategy includes all participants in their originally assigned groups and uses the observed outcome value, regardless of whether they completed the treatment or adhered to the study protocol. This strategy, also known as the treatment policy strategy in ICH E9 , 2019 (R1), is relevant for policy decisions and widely accepted. However, it fails to address ICEs that are terminal events, such as death, since these events directly impact the existence of the outcome variable.
To address the limitation of ITT analyses and approach the issue from a different perspective, the composite variable strategy modifies the causal parameter of interest from the commonly used average causal effect to a comparison between two composite potential outcomes across the treatment arms. The composite potential outcome is defined as the endpoint of interest if no ICE occurs before its measurement, and a pre-specified, clinically meaningful value indicates treatment failure otherwise. The composite variable strategy is widely used in practice across different types of endpoints: binary endpoints (e.g., the non-responder imputation approach), ordinal or continuous endpoints (e.g., Rosenbaum,, 2006), and time-to-event endpoints (e.g., progression-free survival). In addition to the treatment policy strategy and the composite outcome strategy, three additional strategies for addressing ICEs were provided in ICH E9 , 2019 (R1). The hypothetical strategy considers treatment effects under a counterfactual scenario where ICEs do not occur or occur differently. The remaining two strategies–the while-on-treatment strategy and the principal stratification strategy–are less commonly used in practice. ICH E9 , 2019 (R1) recommends using different strategies based on the specific type of ICEs involved. However, very few papers have formally explored how to combine different strategies and rigorously addressed the pitfalls of doing so naively (Lipkovich et al.,, 2020; Qu et al.,, 2021).
1.1 A phase-3 immunology trial
To illustrate our setup and notation, we first introduce a real phase-3 immunology trial (Morand et al.,, 2023; Petri et al.,, 2023), which we will revisit in Section 6. This trial investigates the causal effect of baricitinib versus placebo on Systemic Lupus Erythematosus. Patients in the RCT were assigned either to a baricitinib treatment group or to a placebo control group.
The primary endpoint of interest is an immune response index measured 52 weeks after treatment initiation. Ideally, we would compare this outcome across the two groups at week 52. However, 429 out of 1535 (27.95%) patients experienced ICEs between treatment initiation and the outcome measurement, leading to missingness of the outcome. Figure 1 shows the types and proportions of these ICEs. During the 52 weeks, various ICEs happened. For instance, some patients discontinued the assigned treatment due to adverse events or lack of efficacy, some had protocol deviation and were excluded from the study, some of them withdrew from the study for unknown reasons, and some were lost to follow-up.
Figure 1: Pie chart showing the ICE types and proportions
1.2 Our proposal and contribution
We propose to classify ICEs into two broad types: (1) treatment-unrelated ICEs, such as treatment discontinuation due to relocation or COVID-19 lockdown, which are assumed to be independent of treatment efficacy conditional on the observed covariates; and (2) treatment-related ICEs, such as treatment discontinuation due to adverse events or lack of efficacy, use of rescue medication, and terminating events such as death, which are likely correlated with both treatment and outcomes. Figure 1 illustrates these classifications: yellow solid charts represent treatment-related ICEs, the green solid chart represents a treatment-unrelated ICE, and blue stripped charts represent a mix of both types. For instance, patients withdrew from the study for various reasons–some due to relocation or the pandemic, which can be considered as unrelated to the treatment, while others withdrew due to concerns about potential side effects, which are likely to be treatment-related. When patients did not disclose a detailed reason, we adopt a conservative approach and classify their ICEs as treatment-related.
Both types of ICEs may occur between treatment initiation and outcome measurement. Figure 2 depicts the timeline and three scenarios in our motivating example. Before week 52, patient 1 withdrew from the study due to relocation, which is arguably unrelated to the treatment, patient 2 dropped out of the study due to an adverse event, and patient 3 completed the study through the endpoint at week 52, when the outcome of interest was measured.
Figure 2: Illustration of the motivating immunology trial example
Both types of ICEs lead to missing primary endpoints measured at a landmark time because the outcome cannot be observed or is not even well-defined when an ICE occurs. For the first type of ICE, we apply the hypothetical strategy and assume the ICE event happens at random conditional on the observed covariates. It considers the hypothetical scenario in which the ICE did not occur. For the second category, we utilize the composite variable strategy, incorporating the ICE itself as indicative of treatment failure, for which we provide a formal definition and discussion in Section 3.
A main challenge in combining the two strategies and our key insight is that the two types of ICEs can compete with each other, meaning that patients discontinue due to the first reason that occurs. Use study withdrawal due to relocation and adverse events as examples to represent the two types of ICEs, respectively. For example, suppose a patient discontinues due to relocation like patient 1 in Figure 2. In that case, it remains unknown whether an adverse event would have occurred before the final endpoint measurement in the counterfactual scenario had the withdrawal due to relocation ICE not occurred. It is important to carefully deal with the issue of competing ICEs, which, to the best of our knowledge, has never been pointed out in the literature. We consider this feature and provide the nonparametric identification of our causal parameter of interest under two assumptions: (1) treatment assignment satisfies the ignorability and overlap assumptions, and (2) the treatment-unrelated ICE time is independent of the potential outcomes and the potential values of the treatment-related ICE time, given observed covariates. We present two identification formulas relying on different sets of nuisance models. We then propose two basic estimators based on these two identification formulas and a third estimator that combines them in a similar way to the construction of the augmented inverse probability weighting estimator to achieve double robustness.
The combined estimator can be further improved in terms of both robustness and asymptotic efficiency. To construct a more principled estimator that simultaneously achieves robustness and semiparametric efficiency, we derive the efficient influence function (EIF) thus the semiparametric efficiency bound of our causal parameter. We propose an EIF-based estimator and study its asymptotic properties. We show the EIF estimator is doubly robust, meaning its consistency does not require all nuisance parameters to be consistently estimated—only one of the two sets of nuisance models needs to be correctly specified. Moreover, the EIF estimator achieves the semiparametric efficiency bound under certain conditions on nuisance estimation.
The remaining of the paper proceeds as follows. In Section 2, we provide a motivating example, introduce the basic setup of our research question, describe the composite variable strategy, define our causal parameter of interest, and illustrate the identification challenge. In Section 3, we present the key identification assumptions, provide the nonparametric identification of our causal parameter, and construct three basic estimators. In Section 4, we derive the EIF for our causal parameter, propose a doubly robust and asymptotically efficient estimator based on the EIF, and study its asymptotic properties. In Section 5, we run Monte Carlo simulations to evaluate the finite sample performance of our proposed estimation procedures. In Section 6, we implement the estimators to re-analyze two randomized controlled trials that study the causal effect of baricitinib treatment on systemic lupus erythematosus. In Section 7, we discuss two future research directions.
We use the following notation. Let $\lVert r\rVert_{2}=\{\int r(o)^{2}\textnormal{d}P(o)\}^{1/2}$ denote the $L_{2}(P)$ norm where $P(\cdot)$ denotes the distribution of the observed data $O=o$. For the survival and censoring functions, let $\lVert r\rVert_{2}=\{\iint r(t,o)^{2}\textnormal{d}P(o)\textnormal{d}t\}^{1/2}$ denote the $L_{2}(P)$ norm. We write $b_{n}=O_{P}(a_{n})$ if $b_{n}/a_{n}$ is bounded in probability and $b_{n}=o_{P}(a_{n})$ if $b_{n}/a_{n}$ converges to 00 in probability.
2 Basic setup
2.1 Setup and notation
In this subsection, we introduce our formal setup and mathematical notation. Let $A$ denote the binary treatment indicator, with $A=1$ for patients in the treatment group, and $A=0$ for those in the control group. The primary outcome, denoted as $Y$, is measured at a pre-specified time point $k$. In our motivating example, $Y$ is the immune response index and $k$ is 52 weeks. We adopt the potential outcomes framework (Neyman,, 1923; Rubin,, 1974), where $Y(a)$ denotes the potential outcome under treatment assignment $A=a$ for $a=0,1$, and $Y=Y(A)=AY(1)+(1-A)Y(0)$ is the observed outcome.
As discussed, two types of ICEs may occur between the treatment initiation and the measurement of the outcome: treatment-unrelated ICEs, assumed independent of the potential outcomes conditional on covariates, and treatment-related ICEs, which likely correlate with the endpoints. To formally define the two types of ICEs, let $C$ and $T$ denote time to treatment-unrelated and time-to-treatment-related ICEs, respectively. Both are post-treatment variables, thus we use $C(a)$ and $T(a)$ to denote their potential values under treatment $a$, for $a=0,1$, with observed values $C=C(A)=AC(1)+(1-A)C(0)$ and $T=T(A)=AT(1)+(1-A)T(0)$.
2.2 Causal parameter of interest and challenge in identification
To address treatment-related ICEs, we define a composite potential outcome as $Y^{\textup{cp}}(a)=Y(a)1\{T(a)>k\}$ for $a=0,1$ where the superscript stands for “composite” and define our parameter of interest as the mean contrast, i.e.,
$\tau=E\{Y^{\textup{cp}}(1)-Y^{\textup{cp}}(0)\}=E[Y(1)1\{T(1)>k\}]-E[Y(0)1\{T(0)>k\}].$ (1)
This target parameter is widely used for binary outcomes to capture success or failure, where treatment-related ICEs, such as adverse events, are treated as failure (ICH E9 , 2019, R1). In our motivating example, $Y(a)$ measures whether a patient is an immune responder at week 52 after taking the treatment $A=a$, which is unobserved for patients experiencing a treatment-related ICE such as an adverse event or lack of efficacy. Such patients are defined as non-responders under the composite variable strategy.
The composite variable approach extends beyond binary outcomes to accommodate scenarios where a clinically significant value can denote treatment failure. More specifically, for outcomes such as quality of life, chronic pain, physical functioning, and cognitive performance that are measured on the ordinal and continuous scale, if there is a predefined clinically meaningful value $v$ that can signal treatment failure, then a composite outcome under treatment arm $A=a$ may be constructed as $Y(a)1\{T(a)>k\}+v1\{T(a)\leq k\}$, in which case the causal parameter is defined as in (1) with $Y(a)$ replaced by $Y(a)-v$ for $a=0,1$. Rosenbaum, (2006) discussed similar causal parameters when the ICE is death and the outcome of interest is some measure of life quality after a certain period.
Identifying $\tau$ in the presence of both ICE types is challenging. For clarity, in the remainder of this subsection, we focus on withdrawal due to relocation and adverse events to represent the treatment-unrelated and treatment-related ICEs, respectively. The definition of the causal parameter $\tau$ in (1) essentially treats $Y^{\textup{cp}}(a)$ as the relevant outcome, assigning 0 when adverse events occur, i.e., when $T(a)<k$. If adverse events were the only reason for missingness in the outcome, we could identify $\tau$ using the difference in means of the observed composite outcomes $Y^{\textup{cp}}=Y^{\textup{cp}}(A)=Y1(T>k)$, i.e.,
$\tau$ $=$ $E\{Y1(T>k)\mid A=1\}-E\{Y1(T>k)\mid A=0\}.$
However, this formula is infeasible to learn from observed data because treatment-unrelated ICEs, such as withdrawal due to relocation, could happen before treatment-related ICEs thus censoring $T(a)$. Figure 2 shows how withdrawal due to relocation in patient 1 censored their adverse event time and the outcome, creating missingness in the composite outcome $Y1(T>k)$ for such patients. The key challenge is that the two types of ICEs compete with each other, and patients discontinue due to the first ICE that occurs. Table 1 summarizes observed ICE types and composite outcomes, highlighting that censoring complicates the identification of $\tau$. When there is no ICE, we observe the outcome $Y$, which equals the composite outcome because $T>k$. When we observe an adverse event, the composite outcome is by definition 0. When we observe a withdrawal due to relocation, both the outcome of interest and the time-to-adverse event are missing, and so is the composite outcome. We address the challenge and present identification results in the following section.
Table 1: Summary of the observed ICE types and outcome. $C$ is the time to withdrawal due to relocation (WR), $T$ is the time to adverse event (AE), $Y$ is the outcome of interest, and $k$ is the pre-specified time point when the measurement of $Y$ is taken.
observed ICE type $(T,C,k)$-relationship $Y$ $1(T>k)$ composite outcome $Y^{\textup{cp}}$
no AE/WR $C\wedge T>k$ $Y$ 1 $Y$
AE $C\wedge k>T$ 0 0
WR $T\wedge k>C$
3 Identification and three basic estimators
3.1 Assumptions
We assume the joint distribution of $\{X_{i},A_{i},T_{i}(1),T_{i}(0),C_{i}(1),C_{i}(0),Y_{i}(1),Y_{i}(0)\}$ for patient $i$ is independently and identically distributed from a super population and will drop the subscript $i$ whenever there is no confusion.
We first assume the following ignorability and overlap assumption for the identification of our causal parameter of interest.
Assumption 1 (Treatment assignment)
We assume the following conditional independence and overlap conditions for the treatment assignment:
  • (a)
    $A\perp\!\!\!\perp\{Y(a),T(a),C(a)\}\mid X$ for $a=0,1$.
  • (b)
    For some constant $\eta\in(0,0.5)$, $\eta<e(X)<1-\eta$ with probability 1, where $e(X)=\textup{pr}(A=1\mid X)$ denotes the propensity score.
Assumption 1 is guaranteed by the design of an RCT, allowing our results to apply directly to RCTs like our motivating example. However, our formulation is more general as it can also accommodate observational studies. Under Assumption 1, we have the following equations for $\tau$
$\tau$ $=$ $E[E\{Y1(T>k)\mid A=1,X\}-E\{Y1(T>k)\mid A=0,X\}]$
$=$ $E\left\{\frac{AY1(T>k)}{e(X)}-\frac{(1-A)Y1(T>k)}{1-e(X)}\right\},$
which are the classic outcome regression and the inverse probability weighting identification formulas when we treat the composite potential outcomes $Y^{\textup{cp}}(a)=Y(a)1\{T(a)>k\}$ as the “new” outcome variable of interest with $Y^{\textup{cp}}=Y1(T>k)$. However, as discussed in Section 2.2, they do not serve as identification formulas anymore with the presence of the competing treatment-unrelated ICEs because $Y1(T>k)$ is no longer fully observed if treatment-unrelated ICE happened before both $T$ and $k$. Therefore, we need further identification assumptions on the time of treatment-unrelated ICE. Specifically, we assume it is independent of all other potential values, including the potential outcomes, which is equivalent to the censoring at random assumption in the survival analysis literature when we treat treatment-unrelated ICEs as censoring. In addition, we need the censoring probability to be bounded away from 1. We summarize them in the following assumption.
Assumption 2 (Time of treatment-unrelated ICE)
We assume the following conditional independence and positivity conditions for the time of treatment-unrelated ICE.
  • (a)
    $C(a)\perp\!\!\!\perp\{Y(a),T(a)\}\mid X$ for $a=0,1$.
  • (b)
    For some constant $\eta_{C}>0$, $\textup{pr}\{C(a)>k\mid X\}>\eta_{C}$ with probability 1.
In our motivating example, Assumption 2 states that the time to treatment-unrelated ICEs, such as withdrawal due to relocation, is independent of both the potential outcome of whether a patient will be a responder at week 52 and the time to an adverse event, conditional on observed baseline covariates. In other words, after accounting for factors such as region, baseline corticosteroid use, and the physician global assessment score, the decision to withdraw due to relocation does not provide additional information about the patients’ eventual response or the time to adverse events. Additionally, it assumes that every patient has a nonzero probability of not experiencing a treatment-unrelated ICE.
3.2 Two identification formulas
Using the composite variable strategy for the ICEs that are related to the treatment, recall that we define the parameter of interest as:
$\tau=\mu_{1}-\mu_{0}=E[Y(1)1\{T(1)>k\}]-E[Y(0)1\{T(0)>k\}].$
Under Assumptions 1 and 2, we show the parameter of interest $\tau$ is nonparametrically identified and provide two identification formulas in the following theorem.
Theorem 1 (Nonparametric identification of $\tau$)
Under Assumptions 1 and 2, $\tau$ is nonparametrically identified by two distinct formulas. First,
$\tau$ $=$ $E\left\{\mu_{1}(X)S_{1}(k\mid X)-\mu_{0}(X)S_{0}(k\mid X)\right\},$ (2)
where $\mu_{a}(X)=E(Y\mid T\wedge C>k,X,A=a)$ is the conditional mean of observed outcome among those with no ICE and $A=a$ and $S_{a}(t\mid X)=\textup{pr}(T>t\mid X,A=a)$ is the conditional survival function of treatment-related ICE time among those with $A=a$. Second,
$\tau$ $=$ $E\left[\frac{AY1(T\wedge C>k)}{e(X)G_{1}(k\mid X)}-\frac{(1-A)Y1(T\wedge C>k)}{\left\{1-e(X)\right\}G_{0}(k\mid X)}\right],$ (3)
where $e(X)=\textup{pr}(A=1\mid X)$ is the propensity score and $G_{a}(t\mid X)=\textup{pr}(C>t\mid X,A=a)$ for $a=0,1$ is the survival function of $C$ in the subsample $A=a$.
The first formula (2) expresses $\tau$ using a standard outcome regression approach. It models the conditional mean of the outcome among individuals who have not experienced ICE by $k$, adjusting for the probability of this event occurring. Equation (3) provides an alternative identification using inverse probability weighting. It identifies $\tau$ by averaging observed outcomes only among individuals who have not experienced ICE by time $k$ while correcting for the bias due to differences in the probability of observing a composite outcome. The weights adjust for two key factors: $e(X)$ that accounts for difference in treatment assignment probabilities, and $G_{a}(k\mid X)$ that adjusts for the likelihood of censoring in each treatment group. For instance, units with a lower probability of being observed are up-weighted, while those more likely to remain observed are down-weighted.
Both equations (2) and (3) are valid identification formulas because the nuisance parameters involved are either functions of observed data or identified from observed data. More specifically, nuisance parameters $e(X)$ and $\mu_{a}(X)$ are functions of the observed data. The survival functions of the treatment-related time $S_{a}(t\mid X)$ and the censoring mechanism $G_{a}(t\mid X)$ for $t\leq k$, although not direct functions of observed data, are both identifiable due to the classic result on survival data with censoring at random (Robins and Rotnitzky,, 1992; Robins and Finkelstein,, 2000). For example, $S_{1}(t\mid X)=\textup{pr}(T>t\mid X,A=1)$ for $t\leq k$ can be estimated using observations in the treatment subgroup from treating $T\wedge C\wedge k$ as the observed event time and $1(C\wedge k\geq T)$ as the event status for the treatment-related ICE, and estimating using a Cox proportional hazard model (Cox,, 1972). We can use a similar procedure to estimate $G_{1}(t\mid X)$ by switching the roles of $T$ and $C$, and use similar procedures to estimate $S_{0}(t\mid X)$ and $G_{0}(t\mid X)$ by using observations in control subgroup.
3.3 Two basic estimators
Identification formulas (2) and (3) in Theorem 1 motivate two basic estimators based on estimating the corresponding nuisance functions. Let $\hat{S}_{a}(t\mid X)$ and $\hat{G}_{a}(t\mid X)$ denote the fitted models for $t\leq k$ and $a=0,1$, $\hat{\mu}_{a}(X)$ denote the fitted outcome models for $a=0,1$, and $\hat{e}(X)$ denote the fitted propensity score model. Plugging in the fitted values of the nuisance parameters and taking empirical analogs of equations (2) and (3) motivate the following two estimators, respectively,
$\hat{\tau}^{\textup{out}}$ $=$ $n^{-1}\sum_{i=1}^{n}\hat{\mu}_{1}(X_{i})\hat{S}_{1}(k\mid X_{i})-n^{-1}\sum_{i=1}^{n}\hat{\mu}_{0}(X_{i})\hat{S}_{0}(k\mid X_{i}),$
$\hat{\tau}^{\textup{ipw}}$ $=$ $n^{-1}\sum_{i=1}^{n}\frac{A_{i}Y_{i}1(T_{i}\wedge C_{i}>k)}{\hat{e}(X_{i})\hat{G}_{1}(k\mid X_{i})}-n^{-1}\sum_{i=1}^{n}\frac{(1-A_{i})Y_{i}1(T_{i}\wedge C_{i}>k)}{\left\{1-\hat{e}(X_{i})\right\}\hat{G}_{0}(k\mid X_{i})}.$
The outcome regression estimator $\hat{\tau}^{\textup{out}}$ is consistent if the outcome model and the survival function are correctly specified. The weighting estimator $\hat{\tau}^{\textup{ipw}}$ is consistent if the propensity score model and the censoring mechanism are correctly specified.
3.4 Propensity score-augmented weighting estimator
Combining the two identification formulas (2) and (3), we have the following augmented weighting identification formula,
$\tau$ $=$ $E\left\{\frac{AY1(T\wedge C>k)}{e(X)G_{1}(k\mid X)}-\frac{A-e(X)}{e(X)}\mu_{1}(X)S_{1}(k\mid X)\right\}$ (4)
$-E\left[\frac{(1-A)Y1(T\wedge C>k)}{\left\{1-e(X)\right\}G_{0}(k\mid X)}-\frac{e(X)-A}{1-e(X)}\mu_{0}(X)S_{0}(k\mid X)\right].$
Based on (4), we construct an estimator that augments the weighting estimator $\hat{\tau}^{\textup{ipw}}$ using the estimated outcome models,
$\hat{\tau}^{\textup{aug}}$ $=$ $\hat{\tau}^{\textup{ipw}}-n^{-1}\sum_{i=1}^{n}\left\{\frac{A_{i}-\hat{e}(X_{i})}{\hat{e}(X_{i})}\hat{\mu}_{1}(X_{i})\hat{S}_{1}(k\mid X_{i})+\frac{A_{i}-\hat{e}(X_{i})}{1-\hat{e}(X_{i})}\hat{\mu}_{0}(X_{i})\hat{S}_{0}(k\mid X_{i})\right\}.$
It has a similar mathematical form to the classic augmented inverse probability weighting estimator for the average treatment effect (Bang and Robins,, 2005). We provide the properties of $\hat{\tau}^{\textup{aug}}$ in the following Proposition 1.
Proposition 1 (Double robustness of $\hat{\tau}^{\textup{aug}}$)
Suppose Assumptions 1 and 2 hold, and assume that $G_{a}(k\mid X)$ is correctly specified for $a=0,1$. $\hat{\tau}^{\textup{aug}}$ is a consistent estimator for $\tau$ if either the propensity score model $e(X)$ is correct, or both the outcome regression $\mu_{a}(X)$ and the survival models $S_{a}(k\mid X)$ are correct for $a=0,1$.
Proposition 1 shows that the consistency of $\hat{\tau}^{\textup{aug}}$ depends on the correct model of $G_{a}(k\mid X)$ for $a=0,1$ but remains robust to misspecification of the propensity score model $e(X)$. To build intuition, Equation (4) can be viewed as a projection of the weighting identification formula (3) onto the nuisance tangent space of the propensity score model $A\mid X$, thus we refer to $\hat{\tau}^{\textup{aug}}$ as a propensity score-augmented estimator. However, since it is not further projected onto the tangent space of the $G_{a}(k\mid X)$ model, the estimator does not retain robustness to misspecification of $G_{a}(k\mid X)$. We will further address the issue in Section 4, where we introduce an estimator that is robust to the misspecification of both models.
$\hat{\tau}^{\textup{aug}}$ dominates the weighting estimator $\hat{\tau}^{\textup{ipw}}$ in terms of robustness. The consistency of $\hat{\tau}^{\textup{aug}}$ is contingent on less restrictive requirements on nuisance parameter estimation in the sense that $\hat{\tau}^{\textup{aug}}$ is consistent whenever $\hat{\tau}^{\textup{ipw}}$ is. Consistency of $\hat{\tau}^{\textup{aug}}$ requires the correct specification of the censoring mechanism $G_{a}(k\mid X)$ for $a=0,1$. Given that $G_{a}(k\mid X)$ is correctly modeled, $\hat{\tau}^{\textup{aug}}$ is doubly robust because it is consistent if either the propensity score model is correct, or the outcome regression and the survival models are both correct. There is no clear dominance between $\hat{\tau}^{\textup{aug}}$ and $\hat{\tau}^{\textup{out}}$ as they require consistent estimation of different sets of nuisance parameters. For all three estimators $\hat{\tau}^{\textup{out}}$, $\hat{\tau}^{\textup{ipw}}$, and $\hat{\tau}^{\textup{aug}}$, we can construct variance estimators using nonparametric bootstrap.
4 Another doubly robust and semiparametrically efficient estimator
The propensity score-augmented weighting estimator $\hat{\tau}^{\textup{aug}}$ improves robustness. However, it is not robust to the misspecification of the censoring mechanism, nor does it achieve the semiparametric efficiency bound. In this section, we show that $\hat{\tau}^{\textup{aug}}$ can be further improved. We derive the semiparametric efficient influence function (EIF) for $\tau$ to learn the best asymptotic efficiency a consistent estimator of $\tau$ can achieve and propose an asymptotically efficient and doubly robust estimator based on the EIF.
We first describe the full and observed data structures in our setting and introduce some additional notation. Ideally, we want to observe the full data $(X,Y^{\textup{cp}}(1),Y^{\textup{cp}}(0))$. The missingness of the full data comes from two strings. First, for a given treatment arm $a=0,1$, the treatment-unrelated ICE time $C(a)$ censors the full data because $Y^{\textup{cp}}(a)$ is only observable if $C(a)\geq\{T(a)\wedge k\}$. Within each treatment group, we do not observe the full data due to censoring and only observe $(\Delta(a)=1\{C(a)\geq T(a)\wedge k\},\tilde{T}(a)=C(a)\wedge T(a)\wedge k,\Delta(a)Y^{\textup{cp}}(a))$. Second, the treatment assignment generates another level of missingness, because, for each observation, we never simultaneously observe both composite potential outcomes $\{Y^{\textup{cp}}(1),Y^{\textup{cp}}(0)\}$ even without censoring. Therefore, the observed data is $O=(A,\Delta=\Delta(A),\tilde{T}=\tilde{T}(A),\Delta Y^{\textup{cp}}=\Delta(A)Y^{\textup{cp}}(A))$.
4.1 EIF and estimator
In the following theorem, we provide the EIF for $\tau$.
Theorem 2 (EIF for $\tau$)
Under the nonparametric model, the EIF for $\mu_{1}$ is
$D_{1}(O)$ $=$ $\frac{A}{e(X)}\left\{\frac{Y1(T\wedge C>k)}{G_{1}(k\mid X)}+\mu_{1}(X)S_{1}(k\mid X)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G_{1}}(t)}{S_{1}(t\mid X)G_{1}(t\mid X)}\right\}$ (5)
$-\frac{A-e(X)}{e(X)}\mu_{1}(X)S_{1}(k\mid X)-\mu_{1},$
the EIF for $\mu_{0}$ is
$D_{0}(O)$ $=$ $\frac{1-A}{1-e(X)}\left\{\frac{Y1(T\wedge C>k)}{G_{0}(k\mid X)}+\mu_{0}(X)S_{0}(k\mid X)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G_{0}}(t)}{S_{0}(t\mid X)G_{0}(t\mid X)}\right\}$ (6)
$-\frac{e(X)-A}{1-e(X)}\mu_{0}(X)S_{0}(k\mid X)-\mu_{0},$
and thus, the EIF for $\tau$ is $D_{\tau}(O)=D_{1}(O)-D_{0}(O)$, where $\textnormal{d}M_{G_{a}}(t)=1(C\in\textnormal{d}t,\Delta=0)-1(\tilde{T}\geq t)\textnormal{d}\Lambda_{a}(t\mid X)$ with $\Lambda_{a}(t\mid X)$ denoting the conditional cumulative hazard function for the censoring $C$ in the treatment group $A=a$ for $a=0,1$.
The $M_{G_{a}}(t)$ in the EIF is the martingale constructed from the censoring counting process. Intuitively, $1(C\in\textnormal{d}t,\Delta=0)$ is the actual observed increment in the censoring counting process at time $t$, which records whether a censoring event has occurred, while $1(\tilde{T}\geq t)\textnormal{d}\Lambda_{a}(t\mid X)$ represents the expected increment in the counting process, given the history up to time $t$. The martingale $M_{G_{a}}(t)$ is essentially capturing the difference between the actual observed events and their expected occurrences. The EIF implies another identification formula for $\tau$ by the property that $E\{D_{\tau}(O)\}=0$. Rearranging terms, we have
$\tau$ $=$ $E\left\{\frac{AY1(T\wedge C>k)}{e(X)G_{1}(k\mid X)}-\frac{A-e(X)}{e(X)}\mu_{1}(X)S_{1}(k\mid X)\right\}$ (7)
$-E\left[\frac{(1-A)Y1(T\wedge C>k)}{\left\{1-e(X)\right\}G_{0}(k\mid X)}-\frac{e(X)-A}{1-e(X)}\mu_{0}(X)S_{0}(k\mid X)\right]$
$+E\left[\frac{A}{e(X)}\mu_{1}(X)S_{1}(k\mid X)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G_{1}}(t)}{S_{1}(t\mid X)G_{1}(t\mid X)}\right]$
$-E\left[\frac{1-A}{1-e(X)}\mu_{0}(X)S_{0}(k\mid X)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G_{0}}(t)}{S_{0}(t\mid X)G_{0}(t\mid X)}\right],$
where the first two lines are the same as in the propensity score-augmented inverse probability weighting identification formula (4). Intuitively, we further augment (4) by the term (7) to achieve robustness to misspecification of the censoring model. The EIFs in Theorem 2 motivate the following estimators for $\tau$,
$\hat{\tau}^{\textup{eif}}$ $=$ $\hat{\tau}^{\textup{aug}}+n^{-1}\sum_{i=1}^{n}\frac{A_{i}}{\hat{e}(X_{i})}\hat{\mu}_{1}(X_{i})\hat{S}_{1}(k\mid X_{i})\int_{0}^{\tilde{T}_{i}}\frac{\textnormal{d}M_{\hat{G}_{1}}(t)}{\hat{S}_{1}(t\mid X_{i})\hat{G}_{1}(t\mid X_{i})}$
$-n^{-1}\sum_{i=1}^{n}\frac{1-A_{i}}{1-\hat{e}(X_{i})}\hat{\mu}_{0}(X_{i})\hat{S}_{0}(k\mid X_{i})\int_{0}^{\tilde{T}_{i}}\frac{\textnormal{d}M_{\hat{G}_{0}}(t)}{\hat{S}_{0}(t\mid X_{i})\hat{G}_{0}(t\mid X_{i})}.$
More explicitly, taking the first modification term in $\hat{\tau}^{\textup{eif}}$ as an example, because the observed time points are discrete, the estimator of the integration part for observation $i$ is
$\int_{0}^{\tilde{T}_{i}}\frac{\textnormal{d}M_{\hat{G}_{1}}(t)}{\hat{S}_{1}(t\mid X_{i})\hat{G}_{1}(t\mid X_{i})}$ $=$ $\sum_{t\leq\tilde{T}_{i}}\frac{1(\Delta_{i}=0,C_{i}=t)-\hat{\lambda}_{1}(t\mid X_{i})}{\hat{S}_{1}(t\mid X_{i})\hat{G}_{1}(t\mid X_{i})}$ (8)
$=$ $-\sum_{t\leq\tilde{T}_{i}}\frac{\hat{\lambda}_{1}(t\mid X_{i})}{\hat{S}_{1}(t\mid X_{i})\hat{G}_{1}(t\mid X_{i})}+\frac{1(\Delta_{i}=0)}{\hat{S}_{1}(\tilde{T}_{i}\mid X_{i})\hat{G}_{1}(\tilde{T}_{i}\mid X_{i})},$
where $\hat{\lambda}_{1}(t\mid X_{i})$ denotes the estimator of the conditional hazard function of censoring $\lambda_{1}(t\mid X)$. The first term in (8) is a summation of the ratio $-\hat{\lambda}_{1}(t\mid X_{i})/\{\hat{S}_{1}(t\mid X_{i})\hat{G}_{1}(t\mid X_{i})\}$, where the summation is over all observed event time points before $\tilde{T}_{i}$. The second term is 0 for observations that are not right-censored by the treatment-unrelated ICE, and is $1/\{\hat{S}_{1}(\tilde{T}_{i}\mid X_{i})\hat{G}_{1}(\tilde{T}_{i}\mid X_{i})\}$ for observations with a treatment-unrelated ICE happened at time $\tilde{T}_{i}$.
To construct the estimator $\hat{\tau}^{\textup{eif}}$, we need to estimate the following nuisance parameters: the propensity score model $e(X)$, the outcome model $\mu_{a}(X)$ for $a=0,1$, the survival function $S_{a}(t\mid X)$, and the censoring mechanism $G_{a}(t\mid X)$ for $t\leq k$ and $a=0,1$. Nonetheless, the consistency of $\hat{\tau}^{\textup{eif}}$ does not require the correct specification of all four nuisance parameters, as shown in Theorem 3.
4.2 Asymptotic properties of the EIF-based estimator
We next discuss the asymptotic properties of $\hat{\tau}^{\textup{eif}}$. We first introduce additional notation. For $t\leq k$ and $a=0,1$, let $e^{*}$, $G_{a}^{*}$, $\mu_{a}^{*}$, and $S_{a}^{*}$ denote the probability limit of the estimated propensity score, estimated censoring mechanism, estimated outcome model, and estimated survival function of the AE time, respectively, i.e., $\lVert\hat{e}-e^{*}\rVert=o_{P}(1)$, $\lVert\hat{G}_{a}-G_{a}^{*}\rVert=o_{P}(1)$, $\lVert\hat{\mu}_{a}-\mu_{a}^{*}\rVert=o_{P}(1)$, and $\lVert\hat{S}_{a}-S_{a}^{*}\rVert=o_{P}(1)$. If the propensity score model is consistently estimated, we have $e^{*}=e$, and analogous results for the other three nuisance models.
In the following theorem, we provide the double robustness of the EIF estimator $\hat{\tau}^{\textup{eif}}$.
Theorem 3 (Double robustness of $\hat{\tau}^{\textup{eif}}$)
Under Assumptions 1 and 2, $\hat{\tau}^{\textup{eif}}$ is doubly robust in the sense that it is consistent for $\tau$ if either $\{e^{*}(X)=e(X),G_{a}^{*}(t\mid X)=G_{a}(t\mid X)\}$ or $\{\mu_{a}^{*}(X)=\mu_{a}(X),S_{a}^{*}(t\mid X)=S_{a}(t\mid X)\}$ for $t\leq k$ and $a=0,1$.
Theorem 3 shows that $\hat{\tau}^{\textup{eif}}$ is consistent if at least one of the following two sets of nuisance parameters are consistently estimated: (1) the outcome model $\mu_{a}(X)$ and the survival function $S_{a}(t\mid X)$; (2) the propensity score model $e(X)$ and the censoring mechanism $G_{a}(t\mid X)$, for $t\leq k$ and $a=0,1$. When our data is from an RCT, i.e., the propensity score $e(X)$ is known, and thus $\hat{e}(X)$ is always correctly specified, the EIF estimator $\hat{\tau}^{\textup{eif}}$ is a consistent estimator for $\tau$ if either the censoring mechanism $G_{a}(t\mid X)$ is consistently estimated, or both the outcome model $\mu_{a}(X)$ and the survival function $S_{a}(t\mid X)$ are consistently estimated, for $t\leq k$ and $a=0,1$.
To conduct statistical inference, we provide the asymptotic distribution of $\hat{\tau}^{\textup{eif}}$. We first introduce three technical conditions and then state the asymptotic result in a theorem.
Assumption 3 (Consistency of the nuisance parameters)
Assume that either $\{e^{*}(X)=e(X),\allowbreak G_{a}^{*}(t\mid X)=G_{a}(t\mid X)\}$ or $\{\mu_{a}^{*}(X)=\mu_{a}(X),S_{a}^{*}(t\mid X)=S_{a}(t\mid X)\}$ for $t\leq k$ and $a=0,1$ is satisfied.
Assumption 4 (Donsker Condition)
The class of functions $\{(e,G_{a},\mu_{a},S_{a}):||e-e^{*}||<\delta,||G_{a}-G_{a}^{*}||<\delta,||\mu_{a}-\mu_{a}^{*}||<\delta,||S_{a}-S_{a}^{*}||<\delta\}$ is Donsker for some $\delta>0$.
Assumption 5 (Convergence rates of nuisance parameters)
The convergence rate of the nuisance parameters estimation satisfy $\{||\hat{e}-e^{*}||+||\hat{G}_{a}-G_{a}^{*}||\}\{||\hat{\mu}_{a}-\mu_{a}^{*}||+||\hat{S}_{a}-S_{a}^{*}||\}=o_{P}(n^{-1/2})$ for $a=0,1$.
Assumption 3 requires either the propensity score and censoring models or the survival and outcome models to be consistently estimated. It guarantees the consistency of $\hat{\tau}^{\textup{eif}}$. Assumption 4 imposes restrictions on the nuisance model complexity and is a standard regularity condition (Van der Vaart,, 2000). Recent literature documents that we can use cross-fitting in the estimation of nuisance parameters to relax the Donsker condition (van der Laan et al.,, 2011; Chernozhukov et al.,, 2018; Díaz,, 2019). Assumption 5 imposes additional restrictions on the rate of convergence of the nuisance parameters, in addition to all of them being consistently estimated.
Theorem 4 (Asymptotic distribution)
Under Assumptions 1– 5, the EIF-based estimator satisfies
$n^{1/2}(\hat{\tau}^{\textup{eif}}-\tau)\ =\ n^{-1/2}\sum_{i=1}^{n}D_{\tau}(O_{i})+o_{P}(1).$
$\hat{\tau}^{\textup{eif}}$ is a consistent and asymptotically Gaussian estimator of $\tau$ with asymptotic variance equal to $E\{D_{\tau}^{2}(O)\}$, thus achieves the semiparametric efficiency bound.
Therefore, we can construct the variance estimator based on the semiparametric efficiency bound by taking the empirical analog of the plug-in estimation $n^{-1}\sum_{i=1}^{n}\hat{D}_{\tau}^{2}(O_{i})$. We can also use a nonparametric bootstrap to estimate variance and conduct statistical inference if the nuisance estimators satisfy certain smoothness conditions.
Compare $\hat{\tau}^{\textup{eif}}$ with the propensity score-augmented weighting estimator $\hat{\tau}^{\textup{aug}}$. On the one hand, $\hat{\tau}^{\textup{eif}}$ is more robust in the sense that it allows for misspecification of the censoring model and is asymptotically more efficient. Continuing the intuition discussed in Section 3.4, $\hat{\tau}^{\textup{aug}}$ only orthogonalizes the part corresponding to $A\mid X$ but not $C\mid X,A$, making it robust only to misspecification of the $A\mid X$ model, not the $C\mid X,A$ model. In contrast, $\hat{\tau}^{\textup{eif}}$ accounts for both possible sources of misspecification, leading to greater robustness and efficiency. On the other hand, $\hat{\tau}^{\textup{aug}}$ is more straightforward to implement in practice, as it only requires the estimation of $S_{a}(k\mid X)$ and $G_{a}(k\mid X)$ at time $k$ for $a=0,1$, while $\hat{\tau}^{\textup{eif}}$ requires consistent estimation of the entire survival curves $S_{a}(t\mid X)$ and $G_{a}(t\mid X)$ for $t\leq k$ and $a=0,1$. We consider the choice between $\hat{\tau}^{\textup{eif}}$ and $\hat{\tau}^{\textup{aug}}$ as a trade-off between implementation simplicity and robustness and efficiency of the estimators.
5 Simulation
In this section, we conduct Monte Carlo simulations to study the finite sample performance of our proposed estimators. We first generate the covariates $X=(X_{1},X_{2},X_{3})^{\tiny\textsc{t}}\in\mathbb{R}^{3}$ from three independent standard Gaussian distributions and denote $\tilde{X}_{j}=\{(X_{j}+2)^{2}-1\}/\sqrt{12}$ for $j=1,2,3$. Next, we generate the treatment assignment following $A\mid X\sim\textup{Bernoulli}(\{e(X)\})$ with $e(X)$ being the propensity score model, and generate the potential outcomes following $Y(a)\mid X\sim\mathcal{N}(\mu_{a}(X),\sigma_{a}^{2})$ with $\mu_{a}(X)$ being the outcome model for $a=0,1$. We then generate the potential values of the treatment-related ICE survival time $T(a)$ from the distribution with a survival function $S_{a}(t\mid X)=\exp[-0.002t^{1.2}\exp\{\gamma_{a}(X)\}]$ and the potential values of the treatment-unrelated ICE survival time $C(a)$ from the distributions with a survival function $G_{a}(t\mid X)=\exp[-\varrho_{a}(t)\exp\{\delta_{a}(X)\}]$, for $a=0,1$. For each of the nuisance parameters, we consider two scenarios when it will be correctly modeled and misspecified, therefore, we generate data following two different choices of each. We summarize the detailed data-generating choices in Table 2.
Table 2: Model choices for the nuisance parameters
Correctly specified Misspecified
$\textup{logit}\{e(X)\}$ $(X_{1}+X_{2}+X_{3})/5$ $1(X_{1}\geq 0)\{\exp(\tilde{X}_{2})-X_{2}(1+\tilde{X}_{3})\}-\exp(\tilde{X}_{2})$
$\mu_{1}(X)$ $2(X_{1}+X_{2}+X_{3})$ $1(X_{1}\geq 0)\{X_{2}+\exp(X_{2})\tilde{X}_{3}-\tilde{X}_{2}\}+\tilde{X}_{2}$
$\mu_{0}(X)$ $X_{1}+X_{2}+X_{3}$ $-\tilde{X}_{1}-1(X_{1}>0.5)\tilde{X}_{2}+1(X_{1}<-0.5)X_{2}^{2}\log(|X_{3}|+1)$
$\sigma_{a}$ $0.1(a+1)$ $1$
$\gamma_{1}(X)$ $0.1(X_{1}+2X_{2}-2X_{3})$ $0.1(X_{1}^{2}X_{2}-X_{2}-1)+1(X_{3}\neq 0)X_{2}\log(10X_{3}^{2})$
$\gamma_{0}(X)$ $0.1(X_{1}-2X_{2}+2X_{3})$ $0.01(-\tilde{X}_{1}+\tilde{X}_{2}+\tilde{X}_{3})$
$\delta_{1}(X)$ 00 $0.1(X_{1}^{2}X_{2}-X_{2}-1)+1(X_{3}\neq 0)X_{2}\log(10X_{3}^{2})$
$\delta_{0}(X)$ 00 00
$\varrho_{a}(t)$ $-0.01t^{1.2}$ $-0.01at^{1.2}+0.6*0.01^{1/1.2}(1-a)t$
We consider five data-generating regimes: all four models are correctly specified, misspecified $e$ and correct $(G,\mu,S)$, misspecified $(e,G)$ and correct $(\mu,S)$, misspecified $(\mu,S)$ and correct $(e,G)$, and all four models are misspecified. Let $Y=AY(1)+(1-A)Y(0)$, $T=AT(1)+(1-A)T(0)$, and $C=AC(1)+(1-A)C(0)$. Generate the observed event time as $T\wedge C\wedge k$, the observed event type indicator, and the observed outcome $Y$ if $T\wedge C>k$.
We compare the finite sample performance of the outcome estimator $\hat{\tau}^{\textup{out}}$, the weighting estimator $\hat{\tau}^{\textup{ipw}}$, the augmented weighting estimator $\hat{\tau}^{\textup{aug}}$, and the EIF estimator $\hat{\tau}^{\textup{eif}}$ based on $1000$ Monte Carlo samples with a sample size of $n=1000$. When fitting the nuisance models, we use logistic regression of $A$ on $X$ to estimate the propensity score model, the linear regression of $Y$ on $X$ on the subsample $A=a,T\wedge C>k$ to estimate the outcome model, and the Cox proportional hazard regression of $T$ and $C$ on $X$ on the subsample $A=a$ to estimate the treatment-related ICE survival model and the censoring model, respectively, for $a=0,1$. We use $X$ in all regression models, therefore, when the true data-generating processes involve non-linear functions of $X$, the fitted nuisance models suffer from misspecification. We evaluate the performance of the four estimators with reported bias, standard deviation, and coverage probability using a nonparametric bootstrap with $200$ bootstrap samples in Table 3.
Table 3: Finite sample performance of the four estimators. For each estimator, we report the finite-sample bias, standard deviation (SD), and the coverage rate (CR) of a 95% confidence interval, which is constructed using a nonparametric bootstrap with 200 bootstrap iterations. Each row corresponds to a different data-generating regime.
$\hat{\tau}^{\textup{out}}$ $\hat{\tau}^{\textup{ipw}}$ $\hat{\tau}^{\textup{aug}}$ $\hat{\tau}^{\textup{eif}}$
Bias SD CR Bias SD CR Bias SD CR Bias SD CR
all_correct 0.003 0.100 0.989 0.005 0.154 0.977 0.004 0.152 0.980 0.004 0.110 0.982
$e$_wrong $-0.002$ 0.134 0.978 $-0.338$ 0.527 0.916 $-0.003$ 0.479 0.961 0.000 0.234 0.979
$e\_G$_wrong $-0.002$ 0.130 0.985 $-0.194$ 1.591 0.831 0.141 1.559 0.972 $-0.013$ 0.354 0.969
$\mu\_S$_wrong $-0.085$ 0.170 0.936 $-0.009$ 0.186 0.974 $-0.009$ 0.186 0.974 $-0.012$ 0.176 0.972
all_wrong 0.230 0.233 0.837 0.660 5.799 0.898 0.850 5.782 0.967 0.573 3.022 0.919
Consistent with our theoretical results, all estimators have a finite sample bias close to zero when all models are correctly specified. When $e$ is misspecified and $(G,\mu,S)$ are correct, the weighting estimator $\hat{\tau}^{\textup{ipw}}$ is inconsistent while the other three estimators have small finite sample bias. Under the regime when $e$ and $G$ are misspecified, both $\hat{\tau}^{\textup{ipw}}$ and $\hat{\tau}^{\textup{aug}}$ perform poorly as they have larger finite sample biases, $\hat{\tau}^{\textup{out}}$ and $\hat{\tau}^{\textup{eif}}$ show small finite sample biases as expected. When $(\mu,S)$ are misspecified, the outcome estimator $\hat{\tau}^{\textup{out}}$ has a relatively large bias, while all other three estimators have near-zero biases. All estimators have non-negligible finite sample bias when all nuisance models are misspecified. $\hat{\tau}^{\textup{eif}}$ is most robust to model misspecifications across all different regimes.
$\hat{\tau}^{\textup{eif}}$ shows a relatively smaller standard deviation compared to $\hat{\tau}^{\textup{ipw}}$ and $\hat{\tau}^{\textup{aug}}$ in the regimes where all estimators are consistent. The outcome estimator $\hat{\tau}^{\textup{out}}$ always has the smallest standard deviation when $(\mu,S)$ is correctly specified. Both $\hat{\tau}^{\textup{ipw}}$ and $\hat{\tau}^{\textup{aug}}$ have larger standard deviations across all regimes in which they are consistent. Furthermore, confidence intervals constructed using the nonparametric bootstrap yield valid coverage rates for the corresponding consistent estimators in their respective regimes.
Next, we evaluate two ad-hoc methods commonly used in clinical trials with ICEs, demonstrating through simulations that they are biased in estimating the clinically relevant causal parameter. The first method is the non-responder imputation method that assigns an outcome as 00 whenever an ICE occurs. It is equivalent to using the composite outcome strategy that treats all ICEs as treatment-related, leading to a composite outcome equal to 00. The second method naively applies a hypothetical strategy to all ICEs, assuming they are independent of potential outcomes conditional on observed covariates. We generate data under the “$\mu\_S$_wrong regime, ensuring correct specification of both the propensity score $e(X)$ and the treatment-unrelated ICE survival model $G_{a}(t\mid X)$ for $a=0,1$. Based on $1000$ Monte Carlo simulations with a sample size of $n=1000$, the inverse probability-weighting estimator using the non-responder imputation method exhibits a finite sample bias of $-1.195$ with a standard deviation of $0.077$, and that using the hypothetical strategy for all ICEs has a bias of $1.188$ with a standard deviation of $0.388$. Both ad-hoc methods yield inconsistent estimators.
6 Application
In this section, we re-analyze data from two immunology trials to study the effect of baricitinib on Systemic Lupus Erythematosus (Morand et al.,, 2023; Petri et al.,, 2023). Both trials are double-blinded, randomized, placebo-controlled phase-3 trials. There are three treatment arms: 2mg baritinib, 4mg baritinib, and placebo. The primary endpoint of interest in both studies is the Systemic Lupus Erythematosus Responder Index 4 (SRI4) measured at week 52. It is a binary composite responder index based on both the level of improvement in disease activity and an indicator of no worsening of the overall condition or the development of substantial disease activity in new organ systems. The baseline covariates used in model-fitting include the region each patient is from, the baseline corticosteroid usage, and the baseline value of the Physician Global Assessment score measured at week 0.
There are significant numbers of observations with ICEs happening in both trials (218/760 and 211/775). We classify the observed intercurrent events into two types: those likely to be correlated with the potential outcomes such as adverse events (82.6% and 84.4%), and those conditionally independent of the potential outcomes given covariates and the treatment arm such as withdrawal due to relocation (17.4% and 15.6%). We compare both 2mg and 4mg treatment arms with the placebo group separately. We implemented all four estimators $\hat{\tau}^{\textup{out}}$, $\hat{\tau}^{\textup{ipw}}$, $\hat{\tau}^{\textup{aug}}$, and $\hat{\tau}^{\textup{eif}}$. We used logistic regression for estimating the propensity score and outcome models and used Cox proportional hazards regression to estimate the survival function and censoring model. We used the nonparametric bootstrap with 500 iterations to compute the variance estimator and conduct inference. We summarize the point estimators, their estimated standard errors, and the $p$-values in Table 4.
Table 4: Treatment effect of Baricitinib on the primary endpoint SRI4. The first four columns report the results of Trial 1 and the last four columns report those of Trial 2. The first three rows report the point estimators, standard errors, and the $p$-value of the composite outcome average treatment effect corresponding to the 2mg treatment arm, and the last three rows report those of the 4mg treatment arm. For each trial and each treatment arm, we report the results based on all four estimators.
Trial 1 (Petri et al.,, 2023) Trial 2 (Morand et al.,, 2023)
$\hat{\tau}^{\textup{out}}$ $\hat{\tau}^{\textup{ipw}}$ $\hat{\tau}^{\textup{aug}}$ $\hat{\tau}^{\textup{eif}}$ $\hat{\tau}^{\textup{out}}$ $\hat{\tau}^{\textup{ipw}}$ $\hat{\tau}^{\textup{aug}}$ $\hat{\tau}^{\textup{eif}}$
2mg point 0.030 0.029 0.026 0.026 0.019 0.019 0.022 0.022
se 0.042 0.043 0.042 0.043 0.042 0.043 0.043 0.042
$p$-value 0.479 0.504 0.534 0.540 0.643 0.662 0.602 0.606
4mg point 0.113 0.120 0.115 0.113 $-0.002$ $-0.002$ $-0.002$ $-0.002$
se 0.046 0.046 0.046 0.046 0.042 0.042 0.042 0.042
$p$-value 0.013 0.008 0.012 0.013 0.961 0.962 0.966 0.969
In both trials and both treatment arms, the point estimators are close to each other in most of the cases, and the statistical inference based on different estimators is remarkably robust across methods. Our results suggest a significantly positive effect of 4mg baricitinib treatment on SRI4 in trial 1, although that in trial 2 is insignificant with a slightly negative point estimator. Neither trial observes a significant effect of 2mg baricitinib treatment on SRI4.
Next, we use the two commonly used ad-hoc methods discussed at the end of Section 5 and compare them with our proposed methods. Table 5 reports the estimated treatment effect of both treatment arms across the two trials using the inverse probability-weighting estimator. The non-responder imputation method yields smaller estimates than our proposed methods most of the time, although the difference is modest due to the relatively low proportion of treatment-unrelated ICEs in both trials. Nonetheless, our proposed method targets a clinically more meaningful causal parameter. The second ad-hoc method naively applies the hypothetical strategy to all ICEs, assuming they are independent of potential outcomes conditional on the observed covariates. Such independence assumption is untestable and likely to be violated in clinical trials. For instance, patients discontinued treatment due to adverse events or lack of efficacy. As shown in Table 5, this method produces estimated treatment effects substantially different from the obtained results using our proposed methods, leading to misleading clinical conclusions. It aligns with the large sample bias it exhibits in the simulation study.
Table 5: Treatment effect of Baricitinib on the primary endpoint SRI4 using non-responder imputation (NRI) and hypothetical strategy (HS). The first two columns report the results of Trial 1 and the last two columns report those of Trial 2. The first three rows report the point estimators, standard errors, and the $p$-value of the estimated treatment effect corresponding to the 2mg treatment arm, and the last three rows report those of the 4mg treatment arm. For each trial and each treatment arm, we report the results based on the inverse probability-weighting estimator.
Trial 1 Trial 2
NRI HS NRI HS
2mg point 0.039 0.011 0.010 0.022
se 0.043 0.049 0.042 0.050
$p$-value 0.362 0.815 0.817 0.654
4mg point 0.093 0.073 $-0.008$ 0.044
se 0.044 0.050 0.041 0.051
$p$-value 0.036 0.147 0.847 0.382
7 Discussion
In this study, we address the challenges posed by ICEs in RCTs by proposing methods to handle competing ICEs. We classify ICEs into treatment-related and treatment-unrelated events and apply different strategies for the estimation of a clinically meaningful causal effect. For treatment-related ICEs, we use the composite outcome approach, assigning values that capture treatment failure. For treatment-unrelated ICEs, we adopt a hypothetical strategy that assumes their occurrence is conditionally independent of the potential outcome given covariates. Our key insight is that patients discontinue treatment following the first ICE that occurs, which in turn censors the potential occurrence of ICEs of another type. In this paper, we rigorously address the issue of competing ICEs by employing a formal causal inference framework. We provide identification formulas for the causal parameter of interest, derive its semiparametric efficient influence function, and propose multiple estimators. While our proposed framework provides a rigorous and flexible approach to handling competing ICEs, several challenges and extensions remain open for future research.
7.1 Data collection and ICE classification
Our proposed methods have broad applicability across various therapeutic areas, including immunology, oncology, and cardiology, where treatment discontinuation and other ICEs frequently occur. However, the effectiveness of these approaches relies on the accurate classification of ICEs, which should be performed in collaboration with clinicians to ensure clinical relevance in trial analyses. In our application studies, the two types of ICEs were classified based on a manual review of the detailed comments collected at the clinic. A modernized case report form for a more accurate collection of reasons for treatment discontinuation is required for each application of the proposed method in clinical trials. Fortunately, a cross-industry PHUSE working group is tackling this problem, and a recommended new case report form will be available soon (PHUSE Working Group,, 2024).
7.2 Random $K$
In practice, the measurement of the endpoint of interest $Y$ does not necessarily happen at a fixed point $k$. For instance, the time $K$ that a patient visits the clinic and takes the measurement can be treated as a random variable independent of treatment $A$, potential outcomes $Y(a)$, and both types of ICEs $T(a)$ and $C(a)$ for $a=0,1$. For a random $K$, the analogous formulas to both the previous identification formulas (2) and (3) hold, with a replacement of $k$ to $K$, i.e.,
$\tau$ $=$ $E\left\{\mu_{1}(X)S_{1}(K\mid X)-\mu_{0}(X)S_{0}(K\mid X)\right\}$ (9)
$=$ $E\left[\frac{AY1(T\wedge C>K)}{e(X)G_{1}(K\mid X)}-\frac{(1-A)Y1(T\wedge C>K)}{\left\{1-e(X)\right\}G_{0}(K\mid X)}\right].$ (10)
If $K$ is a pre-treatment covariate that we observe, the previous analyses carry over. If $K$ is also treated as a post-treatment variable, we can construct a weighting estimator following (10). However, (9) ceases to serve as an identification formula unless $S_{a}(K\mid X)$ is identified for $a=0,1$. Considering the presence of three competing time-to-event random variables, it becomes necessary to employ competing risks models. We defer this analysis to future work.
7.3 Comparison between composite variable and principal stratification
We propose to use a combination of the composite variable and hypothetical strategies to deal with ICEs in this project. Following ICH E9 , 2019 (R1), we can also use the combination of principal stratification and hypothetical strategies to deal with the treatment-related and treatment-unrelated ICEs, respectively. Using the principal stratification strategy will target different causal parameters and rely on additional identification assumptions such as monotonicity and principal ignorability of the two potential values of the AE time (Ding and Lu,, 2017; Jiang and Ding,, 2021; Jiang et al.,, 2022). We leave this for future work.
Supplementary Material
We provide proof of all theorems and propositions in the supplementary material.
Appendix A Proofs
A.1 Proof of Theorem 1
First, equation (2) holds because
LHS $=$ $E[E\{Y(1)1(T(1)>k)\mid X\}]$
$=$ $E[E\{Y1(T>k)\mid X,A=1\}]$
$=$ $E\{E(Y\mid T>k,X,A=1)\textup{pr}(T>k\mid X,A=1)\}$
$=$ $E\{E(Y\mid T>k,C>k,X,A=1)\textup{pr}(T>k\mid X,A=1\}=\textup{RHS},$
where the first equality is by the law of iterated expectations, the second equality is by the randomization assumption 1, and the fourth equality is by the censoring at random assumption 2.
Next, equation (3) holds because
RHS $=$ $E\left[E\left\{\frac{AY1(T\wedge C>k)}{e(X)\textup{pr}(C>k\mid X,A=1)}\mid X\right\}\right]$
$=$ $E\left(\frac{1}{e(X)\textup{pr}(C>k\mid X,A=1)}E\left[AY(1)1\{T(1)\wedge C(1)>k\}\mid X\right]\right)$
$=$ $E\left(\frac{1}{e(X)\textup{pr}(C>k\mid X,A=1)}E(A\mid X)E\left[Y(1)1\{T(1)>k\}1\{C(1)>k\}\mid X\right]\right)$
$=$ $E\left(\frac{1}{\textup{pr}(C>k\mid X,A=1)}E[1\{C(1)>k\}\mid X]E\left[Y(1)1\{T(1)>k\}\mid X\right]\right)$
$=$ $E[E\{Y(1)1(T(1)>k)\mid X\}]=\textup{LHS},$
where the first equality is by the law of iterated expectations, the third equality is by the randomization assumption 1, the fourth equality is by the censoring at random assumption 2, and the last equality is again by the law of iterated expectations. $\square$
A.2 Proof of Proposition 1
We prove that when the censoring model is correctly specified, $\hat{\tau}^{\textup{aug}}$ is doubly robust in the sense that it is consistent to $\tau$ if either $e^{*}=e$ or $(\mu_{1}^{*}=\mu_{1},S_{1}^{*}=S_{1})$. Want to show
$E\left\{\frac{AY1(T\wedge C>k)}{e^{*}(X)G_{1}(k\mid X)}-\frac{A-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)-\mu_{1}\right\}$ $=$ 00 0
if either $e^{*}=e$ or $(\mu_{1}^{*}=\mu_{1},S_{1}^{*}=S_{1})$ holds.
$E\left\{\frac{AY1(T\wedge C>k)}{e^{*}(X)G_{1}(k\mid X)}-\frac{A-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right\}$
$=$ $E\left[\frac{E\{AY1(T\wedge C>k)\mid X\}}{e^{*}(X)G_{1}(k\mid X)}-\frac{e(X)-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right]$
$=$ $E\left\{\frac{e(X)}{e^{*}(X)}\mu_{1}(X)S_{1}(k\mid X)-\frac{e(X)-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right\},$
where the first equality is by the law of integrated expectations and the second equality follows from similar derivations as in the proof of Theorem 1. The last line is equal to $\mu_{1}$ if either $e^{*}=e$ or $(\mu_{1}^{*}=\mu_{1},S_{1}^{*}=S_{1})$. $\square$
A.3 Proof of Theorem 2
We follow the semiparametric theory in Bickel et al., (1993) to derive the EIF for $\mu_{1}$. As discussed in Section 4, we have two levels of coarsening of the full data $(X,Y^{\textup{cp}}(1),Y^{\textup{cp}}(0))$, one is because the treatment-unrelated ICE time $C(a)$ is censoring $T(a)\wedge k$, and the other is due to the treatment assignment. Following the steps in Hubbard et al., (2000), we first consider the case when every observation is assigned to the treatment group, i.e., when there is no missingness generated by the treatment assignment, and then project the derived EIF onto the nuisance tangent space of the propensity score to get the final form of EIF for $\mu_{1}$.
With full data $(X,Y^{\textup{cp}}(1),Y^{\textup{cp}}(0))$, the estimating equation for $\mu_{1}$ is $D_{\textup{full}}=Y^{\textup{cp}}(1)-\mu_{1}$. Since $Y^{\textup{cp}}(1)$ is only observable if $C(1)>T(1)\wedge k$, the full data is not available even if all corresponding potential outcomes under treatment $A=1$ are observed, and the observed data is
$O(1)$ $=$ $(X,\Delta(1),\tilde{T}(1),\Delta(1)Y^{\textup{cp}}(1))\sim P_{1},$
where $\Delta(1)=1(C(1)>T(1)\wedge k)$ is the missing indicator with $\Delta(1)=1$ if $Y^{\textup{cp}}(1)$ is observed and $\Delta(1)=0$ otherwise, and $\tilde{T}(1)=T(1)\wedge C(1)\wedge k$ is the observed event time with $\tilde{T}(1)=T(1)\wedge k$ if $\Delta(1)=1$, $Y^{\textup{cp}}(1)$ is observed, and $\tilde{T}(1)=C(1)$ if $\Delta(1)=0$, $Y^{\textup{cp}}(1)$ is missing. An identification formula for $\mu_{1}$ is
$\mu_{1}=E\left\{\frac{\Delta(1)Y^{\textup{cp}}(1)}{G_{1}(\tilde{T}(1)\mid X)}\right\},$
where $G_{1}(t\mid X)=\textup{pr}(C(1)>t\mid X)$ is the probability of not censoring up to time $t$ conditioning on the covariates. We can write the IPCW estimating equation as
$D_{\textup{IPCW}}=\frac{\Delta(1)\{Y^{\textup{cp}}(1)-\mu_{1}\}}{G_{1}(\tilde{T}(1)\mid X)},$
and have $E(D_{\textup{IPCW}})=0$ by the law of iterated expectations and the censoring at random assumption.
Next, we follow the steps in van der Laan and Rubin, (2007) to derive the EIF for $\mu_{1}$ when the observed data is $O(1)=(X,\Delta(1),\tilde{T}(1),\Delta(1)Y^{\textup{cp}}(1))$. For ease of notation, we omit the subscripts, superscripts, or numbers in parenthesis $(1)$ that indicate the potential outcomes under the treatment $A=1$ in the following derivation, so the dependence on the treatment arm is implicit. For ease of notation, denote $T^{1}=T\wedge k$.
Let $T(P)$ denote the tangent space which is the whole Hilbert space $L_{0}^{2}(P)$ since our model is nonparametric. The tangent space can be written as a direct sum of three components $T=T_{X}\oplus T_{\textup{F}}\oplus T_{\textup{CAR}}$ with
$T_{X}$ $=$ $\{h(X)\in L_{0}^{2}(P):E\{h(X)\}=0\},$
$T_{\textup{F}}$ $=$ $\{h(O)\in L_{0}^{2}(P):E\{h(O)\mid C,X\}=0\},$
$T_{\textup{CAR}}$ $=$ $\{h(O)\in L_{0}^{2}(P):E\{h(O)\mid Y^{\textup{cp}},T^{1},X\}=0\},$
where $T_{X}$, $T_{\textup{F}}$, and $T_{\textup{CAR}}$ are orthogonal to each other due to the censoring at random assumption and the factorization of the observed data. These tangent spaces are generated from scores of submodels that perturb the marginal distribution of $X$, the conditional distribution of $Y^{1}\mid X$, and the conditional censoring probability $C\mid X$, respectively. Due to the orthogonality, we have the decomposition of any $h(O)\in L_{0}^{2}(P)$ as
$h(O)$ $=$ $\Pi\{h\mid T(P)\}$
$=$ $\Pi\{h\mid T_{X}(P)\}+\Pi\{h\mid T_{\textup{F}}(P)\}+\Pi\{h\mid T_{\textup{CAR}}(P)\}.$
Let $D(O)$ denote the efficient influence function for $\mu_{1}$. By the results in Chapter 1.4 of van der Laan and Robins, (2003), we have: (1) $D(O)$ should be orthogonal to the nuisance tangent space $T_{\textup{CAR}}$, thus $\Pi\{D(O)\mid T_{\textup{CAR}}(P)\}=0$; (2) $D(O)$ can be rewritten as
$D=D_{\textup{IPCW}}-\Pi\{D_{\textup{IPCW}}\mid T_{\textup{CAR}}(P)\},$
thus
$\Pi\{D\mid T_{\textup{F}}(P)\}$ $=$ $\Pi\{D_{\textup{IPCW}}\mid T_{\textup{F}}(P)\}-\Pi\{\Pi\{D_{\textup{IPCW}}\mid T_{\textup{CAR}}(P)\}\mid T_{\textup{F}}(P)\}$
$=$ $\Pi\{D_{\textup{IPCW}}\mid T_{\textup{F}}(P)\}$
where the last equality is by the fact that $T_{\textup{F}}$ and $T_{\textup{CAR}}$ are orthogonal to each other, and similarly, $\Pi\{D\mid T_{X}(P)\}=\Pi\{D_{\textup{IPCW}}\mid T_{X}(P)\}$.
These projections provide us with two different ways to compute the EIF $D$: (1) directly compute the projections $\Pi\{D_{\textup{IPCW}}\mid T_{X}(P)\}$ and $\Pi\{D_{\textup{IPCW}}\mid T_{\textup{F}}(P)\}$ and sum them up; (2) compute the projection $\Pi\{D_{\textup{IPCW}}\mid T_{\textup{CAR}}(P)\}$ and subtract it from $D_{\textup{IPCW}}$. In the classic survival outcome problem, these two methods are symmetric since $T$ and $C$ are censoring each other thus the projections on $T_{\textup{F}}(P)$ and $T_{\textup{CAR}}(P)$ are very similar. However, due to the complication generated by $Y$ in our setting, the second approach is easier, since the tangent space $T_{\textup{F}}(P)$ is hard to compute.
We next compute $\Pi\{D_{\textup{IPCW}}\mid T_{\textup{CAR}}(P)\}$. By Theorem 1.1 in van der Laan and Robins, (2003),
$T_{\textup{CAR}}(P)=\overline{\left\{\int H(t,\mathcal{F}(t))\textnormal{d}M_{G}(t)\textup{ for all functions }H(t,\mathcal{F}(t))\right\}\cap L_{0}^{2}(P)},$
and the projection of a function $h(O)$ onto $T_{\textup{CAR}}(P)$ is
$\Pi\{h(O)\mid T_{\textup{CAR}}(P)\}=\int_{0}^{\tilde{T}}\left\{E(h(O)\mid dA(t)=1,\mathcal{F}(t))-E(h(O)\mid dA(t)=0,\mathcal{F}(t))\right\}\textnormal{d}M_{G}(t),$
where $A(t)=1(C\leq t)$ is the indicator of whether censoring happens up until time $t$ (define $C=\infty$ if $C>T^{1}$ so that $C$ is always observed), $\mathcal{F}(t)=(\bar{A}(t-),X)$ is the history observed up to time $t$, and $\textnormal{d}M_{G}(t)=1(C\in dt,\Delta=0)-1(\tilde{T}\geq t)d\Lambda(t\mid X)$ is the Doob-Meyer martingale of the counting process of censoring $C$. $E(D_{\textup{IPCW}}\mid dA(t)=1,\mathcal{F}(t))=0$ by the definition of $D_{\textup{IPCW}}$, thus we only need to compute $E(D_{\textup{IPCW}}\mid dA(t)=0,\mathcal{F}(t))$, which plus $\mu_{1}$ is equal to
$E\left\{\frac{\Delta Y^{\textup{cp}}}{G(\tilde{T}\mid X)}\mid dA(t)=0,\mathcal{F}(t)\right\}$
$=$ $E\left\{\frac{\Delta Y^{\textup{cp}}}{G(\tilde{T}\mid X)}\mid T^{1}\geq t,C\geq t,X\right\}$
$=$ $E\left\{\frac{\Delta Y^{\textup{cp}}}{G(\tilde{T}\mid X)}\mid T^{1}\geq t,X\right\}\big{/}G(t\mid X)$
$=$ $E\left\{\frac{Y^{\textup{cp}}}{G(T\wedge k\mid X)}\textup{pr}(C>T\wedge k\mid T^{1}\geq t,T,Y^{1},X)\mid T^{1}\geq t,X\right\}\big{/}G(t\mid X)$
$=$ $E\left\{\frac{Y^{\textup{cp}}}{G(T\wedge k\mid X)}\textup{pr}(C>T\wedge k\mid X)\mid T^{1}\geq t,X\right\}\big{/}G(t\mid X)$
$=$ $E\left\{Y^{\textup{cp}}\mid T^{1}\geq t,X\right\}\big{/}G(t\mid X).$
By the fact that the integration is over $t:t<\tilde{T}$, we have $t<T\wedge k\wedge C$ thus $t>k$. The conditional expectation
$E\left\{Y^{\textup{cp}}\mid T^{1}\geq t,X\right\}$ $=$ $E\left\{Y^{\textup{cp}}\mid T^{1}\geq t,T^{1}=T,X\right\}\textup{pr}(T^{1}=T\mid T^{1}\geq t,X)$
$+E\left\{Y^{\textup{cp}}\mid T^{1}\geq t,T^{1}=k,X\right\}\textup{pr}(T^{1}=k\mid T^{1}\geq t,X)$
$=$ $E(Y\mid T>k\geq t,X)\textup{pr}(T>k\mid T^{1}\geq t,X)$
$=$ $E(Y\mid T>k,X)\textup{pr}(T>k\mid T\wedge k\geq t,X)$
$=$ $E(Y\mid T>k,X)\textup{pr}(T>k\mid T\geq t,X)$
$=$ $\frac{E(Y1(T>k)\mid X)}{\textup{pr}(T>k\mid X)}\frac{\textup{pr}(T>k\mid X)}{\textup{pr}(T\geq t\mid X)}$
$=$ $\frac{E(Y1(T>k)\mid X)}{\textup{pr}(T\geq t\mid X)}.$
Therefore, we have
$\Pi\{D_{\textup{IPCW}}\mid T_{\textup{CAR}}(P)\}+\mu_{1}$ $=$ $-\int_{0}^{\tilde{T}}\frac{E(Y1(T>k)\mid X)}{\textup{pr}(T\geq t\mid X)}\frac{\textnormal{d}M_{G}(t)}{G(t\mid X)}$
$=$ $-E(Y1(T>k)\mid X)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G}(t)}{S(t\mid X)G(t\mid X)},$
and thus the EIF assuming $O(1)$ is the observed data is
$D$ $=$ $\frac{\Delta Y^{\textup{cp}}}{G(\tilde{T}\mid X)}+E(Y^{\textup{cp}}\mid X)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G}(t)}{S(t\mid X)G(t\mid X)}-\mu_{1}$ (S1)
Next, we follow steps in Section 3 of Hubbard et al., (2000) and compute the EIF when the real observed data is $O=(X,A,\Delta,\tilde{T},\Delta Y^{\textup{cp}})$. We need to construct a weighting estimating equation and then subtract its projection onto the nuisance tangent space of the propensity score to get the final form of the EIF. To be clear on the distinction between potential outcomes and the observed values, we add back the dependence on the treatment assignment in (S1) and write it as
$D(1)$ $=$ $\frac{\Delta(1)Y^{\textup{cp}}(1)}{G_{1}(\tilde{T}(1)\mid X)}+E(Y^{\textup{cp}}(1)\mid X)\int_{0}^{\tilde{T}(1)}\frac{\textnormal{d}M_{G_{1}}(t)}{S_{1}(t\mid X)G_{1}(t\mid X)}-\mu_{1}.$
A valid weighting estimating equation is
$D_{\textup{IPW}}$ $=$ $\frac{A}{e(X)}\left\{\frac{\Delta Y^{\textup{cp}}}{G_{1}(\tilde{T}\mid X)}+E(Y^{\textup{cp}}\mid X,A=1)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G_{1}}(t)}{S_{1}(t\mid X)G_{1}(t\mid X)}\right\}-\mu_{1}.$
Further, project this onto the nuisance tangent space of the propensity score, the projection is
$\Pi\{D_{\textup{IPW}}\mid T_{\textup{pscore}}\}$ $=$ $E\left\{D_{\textup{IPW}}\mid A,X\right\}-E\left\{D_{\textup{IPW}}\mid X\right\}$
$=$ $\frac{A-e(X)}{e(X)}E\left\{\frac{\Delta Y^{\textup{cp}}}{G_{1}(\tilde{T}\mid X)}\mid X,A=1\right\}$
$=$ $\frac{A-e(X)}{e(X)}E(Y^{\textup{cp}}\mid X,A=1),$
where the second equality is by the fact that $E\{\textnormal{d}M_{G_{1}}(t)\mid X,A=1\}=E\{\textnormal{d}M_{G_{1}}(t)\mid X\}=0$, and the last equality is by the censoring at random assumption. Therefore, the EIF
$D_{1}$ $=$ $D_{\textup{IPW}}-\Pi\{D_{\textup{IPW}}\mid T_{\textup{pscore}}\}$
$=$ $\frac{A}{e(X)}\left\{\frac{\Delta Y^{\textup{cp}}}{G_{1}(\tilde{T}\mid X)}+E(Y1(T>k)\mid X,A=1)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G_{1}}(t)}{S_{1}(t\mid X)G_{1}(t\mid X)}\right\}$
$-\frac{A-e(X)}{e(X)}E(Y1(T>k)\mid X,A=1)-\mu_{1}$
has the given form in Theorem 2. $\square$
A.4 Proof of Theorem 3
We prove the result by showing
$\mu_{1}=E\left[\frac{A}{e^{*}(X)}\left\{\frac{Y1(T\wedge C>k)}{G_{1}^{*}(k\mid X)}+\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G_{1}^{*}}(t)}{S_{1}^{*}(t\mid X)G_{1}^{*}(t\mid X)}\right\}-\frac{A-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right]$ (S2)
if either $\{e^{*}(X)=e(X),G_{1}^{*}(t\mid X)=G_{1}(t\mid X)\}$ or $\{\mu_{1}^{*}(X)=\mu_{1}(X),S_{1}^{*}(t\mid X)=S_{1}(t\mid X)\}$ for $t\leq k$.
Denote
$\mathcal{T}_{1}$ $=$ $\frac{A}{e^{*}(X)}\frac{Y1(T\wedge C>k)}{G_{1}^{*}(k\mid X)}-\frac{A-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)-\mu_{1}(X)S_{1}(k\mid X),$
$\mathcal{T}_{2}$ $=$ $\frac{A}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G_{1}^{*}}(t)}{S_{1}^{*}(t\mid X)G_{1}^{*}(t\mid X)}.$
By the identification formula (2), to prove (S2), it suffices to show $E(\mathcal{T}_{1}+\mathcal{T}_{2})=0$ if either $\{e^{*}(X)=e(X),G_{1}^{*}(t\mid X)=G_{1}(t\mid X)\}$ or $\{\mu_{1}^{*}(X)=\mu_{1}(X),S_{1}^{*}(t\mid X)=S_{1}(t\mid X)\}$ for $t\leq k$.
First, we have
$E\{\textnormal{d}M_{G_{1}^{*}}(t)\mid X\}$ $=$ $E\{1(C\in\textnormal{d}t,\Delta=0)\mid X\}-E\{1(\tilde{T}\geq t)\textnormal{d}\Lambda_{1}^{*}(t\mid X)\mid X\}$
$=$ $S_{1}(t\mid X)G_{1}(t\mid X)\textnormal{d}\Lambda_{1}(t\mid X)-S_{1}(t\mid X)G_{1}(t\mid X)\textnormal{d}\Lambda_{1}^{*}(t\mid X)$
$=$ $S_{1}(t\mid X)G_{1}(t\mid X)\{\textnormal{d}\Lambda_{1}(t\mid X)-\textnormal{d}\Lambda_{1}^{*}(t\mid X)\},$
therefore,
$E(\mathcal{T}_{2}\mid X)$ $=$ $\frac{e(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)E\left\{\int_{0}^{k}\frac{\textnormal{d}M_{G_{1}^{*}}(t)}{S_{1}^{*}(t\mid X)G_{1}^{*}(t\mid X)}\mid X\right\}$
$=$ $\frac{e(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\int_{0}^{k}\frac{S_{1}(t\mid X)G_{1}(t\mid X)}{S_{1}^{*}(t\mid X)G_{1}^{*}(t\mid X)}\{\textnormal{d}\Lambda_{1}(t\mid X)-\textnormal{d}\Lambda_{1}^{*}(t\mid X)\},$
where the first equality follows from Assumption 1. Next, we have
$E(\mathcal{T}_{1}\mid X)$ $=$ $\frac{e(X)G_{1}(k\mid X)}{e^{*}(X)G_{1}^{*}(k\mid X)}\mu_{1}(X)S_{1}(k\mid X)-\frac{e(X)-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)-\mu_{1}(X)S_{1}(k\mid X)$
$=$ $\frac{e(X)}{e^{*}(X)}\left\{\frac{G_{1}(k\mid X)}{G_{1}^{*}(k\mid X)}-1\right\}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)$
+$+\left\{\frac{e(X)G_{1}(k\mid X)}{e^{*}(X)G_{1}^{*}(k\mid X)}-1\right\}\left\{\mu_{1}(X)S_{1}(k\mid X)-\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right\}$
$=$ $-\frac{e(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\int_{0}^{k}\frac{G_{1}(t\mid X)}{G_{1}^{*}(t\mid X)}\{\textnormal{d}\Lambda_{1}(t\mid X)-\textnormal{d}\Lambda_{1}^{*}(t\mid X)\}$
+$+\left\{\frac{e(X)G_{1}(k\mid X)}{e^{*}(X)G_{1}^{*}(k\mid X)}-1\right\}\left\{\mu_{1}(X)S_{1}(k\mid X)-\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right\},$
where the last equality follows from the Duhamel equation (Gill and Johansen,, 1990; Westling et al.,, 2024). Combing $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, we have
$E(\mathcal{T}_{1}+\mathcal{T}_{2})$ $=$ $E\{E(\mathcal{T}_{1}+\mathcal{T}_{2}\mid X)\}$
$=$ $\frac{e(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\int_{0}^{k}\left\{\frac{S_{1}(t\mid X)}{S_{1}^{*}(t\mid X)}-1\right\}\frac{G_{1}(t\mid X)}{G_{1}^{*}(t\mid X)}\{\textnormal{d}\Lambda_{1}(t\mid X)-\textnormal{d}\Lambda_{1}^{*}(t\mid X)\}$
+$+\left\{\frac{e(X)G_{1}(k\mid X)}{e^{*}(X)G_{1}^{*}(k\mid X)}-1\right\}\left\{\mu_{1}(X)S_{1}(k\mid X)-\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right\},$
which is equal to 0 if either $\{e^{*}(X)=e(X),G_{1}^{*}(t\mid X)=G_{1}(t\mid X)\}$ or $\{\mu_{1}^{*}(X)=\mu_{1}(X),S_{1}^{*}(t\mid X)=S_{1}(t\mid X)\}$ for $t\leq k$.
By symmetry, the part corresponding to the control arm $a=0$ also holds, thus the double robustness in Theorem 3 holds. $\square$
A.5 Proof of Theorem 4
Following the von Mises expansion (Hines et al.,, 2022), let $P$ and $P_{n}$ denote the true and empirical distribution of the observed data, respectively, and $\hat{P}_{n}$ denotes an estimated $P$, we have
$n^{1/2}(\hat{\tau}^{\textup{eif}}-\tau)\ =\ n^{-1/2}\sum_{i=1}^{n}D_{\tau}(O_{i})+n^{1/2}(P_{n}-P)\{D_{\tau}(O,\hat{P}_{n})-D_{\tau}(O)\}-n^{1/2}R(P,\hat{P}_{n}),$
where $D_{\tau}(O,\hat{P}_{n})$ is the EIF for $\tau$ when plugging in the estimated values of the nuisance parameters and $R(P,\hat{P}_{n})$ denotes the higher order remainder term. Under Assumption 4, the empirical process term $n^{1/2}(P_{n}-P)\{D_{\tau}(O,\hat{P}_{n})-D_{\tau}(O)\}=o_{P}(1)$ (Van der Vaart,, 2000). The remainder term corresponding to the treatment arm $a=1$ satisfies
$R_{1}(P,\hat{P}_{n})$ $=$ $-E\{D_{\tau}(O,\hat{P}_{n})\}-E\{\hat{\mu}_{1}(X)\hat{S}_{1}(k\mid X)-\mu_{1}(X)S_{1}(k\mid X)\}$
$=$ $-E\left[\frac{A}{\hat{e}(X)}\left\{\frac{Y1(T\wedge C>k)}{\hat{G}_{1}(k\mid X)}+\hat{\mu}_{1}(X)\hat{S}_{1}(k\mid X)\int_{0}^{k}\frac{\textnormal{d}M_{\hat{G}_{1}}(t)}{\hat{G}_{1}(t\mid X)\hat{S}_{1}(t\mid X)}\right\}\right.$
$\quad\quad\ \left.-\frac{A-\hat{e}(X)}{\hat{e}(X)}\hat{\mu}_{1}(X)\hat{S}_{1}(k\mid X)-\hat{\mu}_{1}(X)\hat{S}_{1}(k\mid X)+\hat{\mu}_{1}(X)\hat{S}_{1}(k\mid X)-\mu_{1}(X)S_{1}(k\mid X)\right]$
$=$ $-E\left[\frac{e(X)}{\hat{e}(X)}\hat{\mu}_{1}(X)\hat{S}_{1}(k\mid X)\int_{0}^{k}\left\{\frac{S_{1}(t\mid X)}{\hat{S}_{1}(t\mid X)}-1\right\}\frac{G_{1}(t\mid X)}{\hat{G}_{1}(t\mid X)}\{\textnormal{d}\Lambda_{1}(t\mid X)-\textnormal{d}\hat{\Lambda}_{1}(t\mid X)\}\right]$
$+E\left[\left\{\frac{e(X)G_{1}(k\mid X)}{\hat{e}(X)\hat{G}_{1}(k\mid X)}-1\right\}\left\{\mu_{1}(X)S_{1}(k\mid X)-\hat{\mu}_{1}(X)\hat{S}_{1}(k\mid X)\right\}\right]$
$=$ $E\left\{\frac{e(X)}{\hat{e}(X)}\hat{\mu}_{1}(X)\hat{S}_{1}(k\mid X)\int_{0}^{k}\frac{S_{1}(t\mid X)-\hat{S}_{1}(t\mid X)}{\hat{S}_{1}(t\mid X)}\frac{G_{1}(t\mid X)-\hat{G}_{1}(t\mid X)}{\hat{G}_{1}(t\mid X)}\textnormal{d}t\right\}$
$+E\left[\frac{e(X)G_{1}(k\mid X)-\hat{e}(X)\hat{G}_{1}(k\mid X)}{\hat{e}(X)\hat{G}_{1}(k\mid X)}\left\{\mu_{1}(X)S_{1}(k\mid X)-\hat{\mu}_{1}(X)\hat{S}_{1}(k\mid X)\right\}\right],$
where the third equality follows from a similar derivation as in the Proof of Theorem 3 and the fourth equality follows from the Duhamel equation (Gill and Johansen,, 1990; Westling et al.,, 2024). Let $R_{11}$ and $R_{12}$ denote the two terms in the last two lines, respectively. For the remainder term to be of small order, we need $R_{11}=o_{P}(n^{-1/2})$ and $R_{12}=o_{P}(n^{-1/2})$.
Next, we show that Assumption 5 is a sufficient condition for $R_{11}=o_{P}(n^{-1/2})$ and $R_{12}=o_{P}(n^{-1/2})$. In the following discussion, we suppress the dependency of the nuisance functions and their estimated values on $X$ for ease of notation. By the facts that $e\hat{\mu}_{1}\hat{S}_{1}(k)/\hat{e}$ is bounded and $1/\{\hat{S}_{1}(t)\hat{G}_{1}(t)\}$ is bounded for any $t\leq k$, we employ the Cauchy-Schwarz inequality to upper bound the term $R_{11}$ by a constant times $\int_{0}^{k}||S_{1}(t)-\hat{S}_{1}(t)||_{2}||G_{1}(t)-\hat{G}_{1}(t)||_{2}\textnormal{d}t$, where $||\cdot||_{2}$ denotes the $L_{2}(P)$ norm. It follows that Assumption 5 guarantees $R_{11}=o_{P}(n^{-1/2})$. For $R_{12}$, we further suppress the dependency of $S_{1}(k\mid X)$, $G_{1}(k\mid X)$ and their estimators on both $k$ and $X$, and rewrite the term as $R_{12}=E[\{\hat{e}(G_{1}-\hat{G}_{1})+G_{1}(e-\hat{e})\}\{\hat{\mu}_{1}(S_{1}-\hat{S}_{1})+S_{1}(\mu_{1}-\hat{\mu}_{1})\}/(\hat{e}\hat{G}_{1})]$. Because $\hat{e}/\hat{G}_{1}$ is bounded, by the Cauchy-Schwarz inequality, the first term in $R_{12}$ is upper bounded by $||G_{1}-\hat{G}_{1}||_{2}||S_{1}-\hat{S}_{1}||_{2}=o_{P}(n^{-1/2})$ by Assumption 5. The other three terms in $R_{12}$ can be similarly bounded. Thus, $R_{1}(P,\hat{P}_{n})=o_{P}(n^{-1/2})$.
Symmetric arguments imply the analogous term for the control arm satisfies $R_{0}(P,\hat{P}_{n})=o_{P}(n^{-1/2})$. Therefore, results in Theorem 4 follow from the central limit theorem. $\square$